# The consistency strength of determinacy when all sets are universally Baire

### v2

@@ 5.4. Preservation of certification under hulls
Let $i<\omega$ and let $\bar{\mathcal{W}}$ be the transitive collapse of a hull of $\mathcal{W}|\Omega$ for some sufficiently large $\Omega$ such that $\mathcal{W}|\delta_{i}\subseteq\bar{\mathcal{W}}$ and $|\bar{\mathcal{W}}|<\kappa_{i}$. Let
$\pi\colon\bar{\mathcal{W}}\rightarrow\mathcal{W}$
be the uncollapse embedding, $\bar{\mathcal{M}}$ the collapse of $\mathcal{M}|\Omega$, and $(\bar{\kappa}_{j},\bar{\delta}_{j}\mid j<\omega)$ the preimages of $(\kappa_{j},\delta_{j}\mid j<\omega)$ (so in particular, $\bar{\kappa}_{j}=\kappa_{j}$ for $j<i$ and $\bar{\delta}_{j}=\delta_{j}$ for $j\leq i$). Then every extender in $\bar{\mathcal{M}}$ with critical point $\bar{\kappa}_{i}$ is certified in $\bar{\mathcal{M}}$.
@@
This lemma can be proved similar as Lemma 5.14, so we only describe the changes that need to be made. It no longer holds that the direct limit $\mathcal{M}_{\infty}^{i,-}$ does not get collapsed, so let ¯$\bar{\mathcal{M}}_{\infty}^{i,-}$ be the preimage of $\mathcal{M}_{\infty}^{i,-}$ under $\pi$. Choose $E,F,\bar{\mathcal{U}},\mathcal{U},\bar{b},b$ as in the proof of Lemma 5.14. As $\bar{\mathcal{U}}$ and $\mathcal{U}$ are iteration trees on different models, i.e., ¯$\bar{\mathcal{M}}_{\infty}^{i,-}$ and $\mathcal{M}_{\infty}^{i,-}$, we cannot simply use hull condensation to argue that $\bar{\mathcal{U}}^{\frown}\bar{b}$ is by $\Sigma^{i,\infty}$ as well. Instead, we consider trees $\bar{\mathcal{T}}$ and $\mathcal{T}$ on $\mathcal{P}_{i}^{-}$ with last models ¯$\bar{\mathcal{M}}_{\infty}^{i,-}$ and $\mathcal{M}_{\infty}^{i,-}$ respectively. Now apply hull condensation to the stacks $\bar{\mathcal{T}}^{\frown}\bar{\mathcal{U}}$ and $\mathcal{T}^{\frown}\mathcal{U}$ to obtain that $\bar{\mathcal{T}}^{\frown}\bar{\mathcal{U}}$ is by $\Sigma_{i}$. This allows us to continue as in the proof of Lemma 5.14 and obtain
$\pi_{E}\upharpoonright\bar{\mathcal{M}}_{\infty}^{i,-}=\pi_{\bar{b}}^{\bar{\mathcal{U}}}=\pi^{\bar{\Sigma}^{i,\infty}}_{\bar{\mathcal{M}}_{\infty}^{i,-},\pi_{E}(\bar{\mathcal{M}}_{\infty}^{i,-})},$
as desired, where $\bar{\Sigma}^{i,\infty}=(\Sigma_{i})_{\bar{\mathcal{M}}_{\infty}^{i,-}}$. ∎
5.5. All extenders on the $\mathcal{M}$-sequence are certified
Similarly as in [13, Lemma 2.25] we have that fully backgrounded extenders with critical point $\kappa_{i}$ for some $i\geq 0$ are certified.
Lemma 5.16.
Let $i\geq 0$ and let $E$ be an extender on $\mathcal{M}$ with critical point $\kappa_{i}$ that is added by the fully backgrounded part of the construction of $\mathcal{M}$. Then $E$ is certified.
Proof.
Let $F$ be an extender with critical point $\kappa_{i}$ that gets added by the fully backgrounded part of the construction of $\mathcal{M}$. Say $\mathcal{N}_{\xi+1}=(J_{\alpha}^{\vec{E}},\in,\vec{E},F)$ and let $F^{*}$ be the background extender of $F$. Let $\mathcal{U}$ be an iteration tree on $\mathcal{M}_{\infty}^{i,-}$ of limit length such that if we let $b=\Sigma^{i,\infty}(\mathcal{U})$,
$\pi^{\Sigma^{i,\infty}}_{\mathcal{M}_{\infty}^{i,-},\pi_{F}(\mathcal{M}_{\infty}^{i,-})}=\pi_{b}^{\mathcal{U}}.$
Moreover, let $\mathcal{U}^{*}$ be an iteration tree on $\mathcal{M}_{\infty}^{i,-}$ of limit length such that if we let $b^{*}=\Sigma^{i,\infty}(\mathcal{U}^{*})$,
$\pi^{\Sigma^{i,\infty}}_{\mathcal{M}_{\infty}^{i,-},\pi_{F^{*}}(\mathcal{M}_{\infty}^{i,-})}=\pi_{b^{*}}^{\mathcal{U}^{*}}.$
@@ 5.5. All extenders on the $\mathcal{M}$-sequence are certified
be the canonical realization map. Then $k\upharpoonright\pi_{F}(\mathcal{M}_{\infty}^{i,-})\colon\pi_{F}(\mathcal{M}_{\infty}^{i,-})\rightarrow\pi_{F^{*}}(\mathcal{M}_{\infty}^{i,-})$ makes the diagram in Figure 2 commute.
$\pi_{F^{*}}$$\pi_{F^{*}}(\mathcal{M}_{\infty}^{i,-})$$\mathcal{M}_{\infty}^{i,-}$$\pi_{F}$$\pi_{F}(\mathcal{M}_{\infty}^{i,-})$$k$
@@
As $F^{*}$ is an extender on $\mathcal{W}$ and $(\Sigma_{i})^{g}$ is strongly guided by term relations for all $g\subseteq\operatorname{Col}(\omega,\mathcal{P}^{i})$,
$\pi_{F^{*}}\upharpoonright\mathcal{M}_{\infty}^{i,-}=\pi_{b^{*}}^{\mathcal{U}^{*}}.$
We now define a cofinal branch $c$ through $\mathcal{U}$ such that $k(\mathcal{M}_{c}^{\mathcal{U}})=\mathcal{M}_{b^{*}}^{\mathcal{U}^{*}}$ and $\pi_{F}\upharpoonright\mathcal{M}_{\infty}^{i,-}=\pi_{c}^{\mathcal{U}}$.
@@
Both $\mathcal{U}^{\frown}c$ and $\mathcal{U}^{*}{}^{\frown}b^{*}$ are iteration trees on $\mathcal{M}_{\infty}^{i,-}$ and $\mathcal{U}^{\frown}c$ is a hull of $\mathcal{U}^{*}{}^{\frown}b^{*}$. As $\mathcal{U}^{*}{}^{\frown}b^{*}$ is according to $\Sigma^{i,\infty}$, by hull condensation $\mathcal{U}^{\frown}c$ is according to $\Sigma^{i,\infty}$ as well and hence
@@
Let $i\geq 0$ and let $E$ be an extender on $\mathcal{M}$ with critical point $\kappa_{i}$ that is added by the generically countably complete part of the construction of $\mathcal{M}$. Then $E$ is certified.
@@
Let $E$ be an extender on the sequence of $\mathcal{M}$ with critical point $\kappa_{i}$ that is added by the generically countably complete part of the construction of $\mathcal{M}$ and suppose toward a contradiction that $E$ is not certified. Let $\hat{\mathcal{W}}$ be a countable elementary substructure of a sufficiently large initial segment $\mathcal{W}|\Omega$ of $\mathcal{W}$. Let
@@
and suppose that all relevant objects are in the range of $\sigma$. For notational simplicity write $\mathcal{M}$ for the result of the $\mathcal{M}$-construction in $\mathcal{W}|\Omega$ and let $\hat{\mathcal{M}}$ be the transitive collapse of $\mathcal{M}$, i.e., the result of the $\mathcal{M}$-construction in $\hat{\mathcal{W}}$. Say $\sigma(\hat{\kappa}_{i},\hat{E},\hat{\mathcal{M}}_{\infty}^{i},\hat{\Sigma}^{i,\infty})=(\kappa_{i},E,\mathcal{M}_{\infty}^{i},\Sigma^{i,\infty})$.
^$\hat{\mathcal{M}}_{\infty}^{i}$ is an element of the direct limit system giving rise to $\mathcal{M}_{\infty}^{i}$. So when comparing ^$\hat{\mathcal{M}}_{\infty}^{i}$ via $\hat{\Sigma}^{i,\infty}$ and $\mathcal{M}_{\infty}^{i}$ via $\Sigma^{i,\infty}$ inside $\mathcal{W}$, ^$\hat{\mathcal{M}}_{\infty}^{i}$ iterates to $\mathcal{M}_{\infty}^{i}$. Moreover,
$\sigma\upharpoonright\hat{\mathcal{M}}_{\infty}^{i,-}=\pi^{\hat{\Sigma}^{i,\infty}}_{\hat{\mathcal{M}}_{\infty}^{i,-},\sigma(\hat{\mathcal{M}}_{\infty}^{i,-})}$
@@
$\pi_{\hat{E}}\upharpoonright\hat{\mathcal{M}}_{\infty}^{i,-}=\pi^{\hat{\Sigma}^{i,\infty}}_{\hat{\mathcal{M}}_{\infty}^{i,-},\pi_{\hat{E}}(\hat{\mathcal{M}}_{\infty}^{i,-})}.$
So $\hat{E}$ is certified via $\hat{\Sigma}^{i,\infty}$. By elementarity of $\sigma$, this implies that $E$ is certified via $\Sigma^{i,\infty}$, a contradiction. ∎
@@ 5.6. Characterization via operators
Now we argue that we can also characterize extenders using operators in order to obtain the extenders in a simply definable way (e.g., in $M_{1}^{\#,\Sigma_{i}}$ of the current level). In what follows, we define universal mice at $\kappa_{i}$ and $\kappa_{i}$-good universal mice as in [13]. The arguments in [13] generalize to $\kappa_{i}$, $i>0$, by considering $\mathcal{M}$-constructions instead of fully backgrounded constructions. In particular, [13, Lemmas 2.15 and 2.16] generalize and we can show that good universal mice exist. We give the details in what follows.
@@
Let $i<\omega$ and let $X\in\mathcal{W}|\delta_{i+1}$ be a premouse such that $\mathcal{M}|\delta_{i}\unlhd X$. We say $\mathcal{Q}\subseteq\mathcal{W}|\delta_{i+1}$ is a universal mouse over $X$ at $\kappa_{i}$ if $\mathcal{Q}$ is an $X$-premouse such that $\mathcal{Q}$ is $(\delta_{i+1},\delta_{i+1})$-iterable with respect to the extenders on its sequence with critical point above $X\cap\operatorname{Ord}$, for every $\eta\in(X\cap\operatorname{Ord},\delta_{i+1})$, $\mathcal{O}_{\eta}^{\mathcal{Q}}\unlhd Lp_{\omega}^{\infty}(\mathcal{Q}|\eta)$ and for a stationary set of $\xi<\delta_{i+1}$, $(\xi^{+})^{\mathcal{Q}}=(\xi^{+})^{\mathcal{W}}$.
Before we can define good universal mice, we need some notation. Let $i<\omega$ and let $X\in\mathcal{W}|\delta_{i+1}$ be a premouse with $X\unrhd\mathcal{M}|\delta_{i}$. For any $X$-premouse $\mathcal{Q}\subseteq\mathcal{W}|\delta_{i+1}$, let $\kappa^{\mathcal{Q}}$ denote the least ${<}\delta_{i+1}$-strong cardinal of $\mathcal{Q}$, if it exists. For any universal mouse $\mathcal{Q}\subseteq\mathcal{W}|\delta_{i+1}$ over $X$ at $\kappa_{i}$, consider, analogous to Section 5.2, the directed system given by
@@
Write $\mathcal{M}_{\infty}^{\mathcal{Q}}$ for the corresponding direct limit and $\delta_{\infty}^{\mathcal{Q}}$ for the largest Woodin cardinal in $\mathcal{M}_{\infty}^{\mathcal{Q}}$. Note that $\mathcal{M}_{\infty}^{\mathcal{Q}}\in\mathcal{Q}$ and let
$\mathcal{R}^{\mathcal{Q}}=\mathcal{M}_{\infty}^{\mathcal{Q}}|((\delta_{\infty}^{\mathcal{Q}})^{+\omega})^{\mathcal{M}_{\infty}^{\mathcal{Q}}}.$
@@
Let $i<\omega$ and let $X\in\mathcal{W}|\delta_{i+1}$ be a premouse such that $\mathcal{M}|\delta_{i}\unlhd X$. We say a universal mouse $\mathcal{Q}$ over $X$ at $\kappa_{i}$ is a $\kappa_{i}$-good universal mouse over $X$ iff it has a $(\delta_{i+1},\delta_{i+1})$-iteration strategy $\Phi$ with respect to extenders with critical point above $X\cap\operatorname{Ord}$ such that whenever $i\colon\mathcal{Q}\rightarrow\mathcal{N}$ is an embedding obtained from an iteration according to $\Phi$, $\mathcal{N}$ is a universal mouse, $\mathcal{R}^{\mathcal{N}}$ is a $\Sigma_{i}$-iterate of $\mathcal{P}^{i}$, and whenever $g$ is generic over $\mathcal{W}$ for a partial order in $\mathcal{N}$,
@@
The existence of good universal mice follows as in [13, Lemmas 2.15 and 2.16], cf. also the proof of Lemma 5.11.
@@
Let $i<\omega$ and let $X\unrhd\mathcal{M}|\delta_{i}$ be a premouse with $X\in\mathcal{W}|\delta_{i+1}$. Fix some ordinal $\eta$ with $X\cap\operatorname{Ord}<\eta<\delta_{i+1}$ and let $\mathcal{Q}=\mathcal{Q}^{X,\eta}$ be the result of an $\mathcal{M}$-construction above $X$ inside $\mathcal{W}|\delta_{i+1}$ where the fully backgrounded part only uses extenders with critical point above $\eta$. Let $\Phi$ be the iteration strategy for $\mathcal{Q}$ induced by the background strategy and let $\mathcal{N}$ be an iterate of $\mathcal{Q}$ (above $X$) according to $\Phi$ such that the iteration embedding $i\colon\mathcal{Q}\rightarrow\mathcal{N}$ exists. Then $\mathcal{N}$ is a $\kappa_{i}$-good universal mouse over $X$.
@@
$\mathcal{R}^{\mathcal{N}}$ is a $\Sigma_{i}$-iterate of $\mathcal{P}^{i}$ and $\mathcal{N}$ is $\kappa_{i}$-good, i.e., whenever $g$ is generic over $\mathcal{W}$ for a partial order in $\mathcal{N}$, $(\Sigma^{g}_{i})_{\mathcal{R}^{\mathcal{N}}}\upharpoonright\mathcal{N}[g]\in L_{1}(\mathcal{N}[g])$, by the argument for Lemma 5.11. So we only sketch the argument that $\mathcal{N}$ is universal and show that for a stationary set of $\xi<\delta_{i+1}$, $(\xi^{+})^{\mathcal{N}}=(\xi^{+})^{\mathcal{W}}$. Let $\mathcal{W}^{*}$ be the iterate of $\mathcal{W}$ obtained by applying the $\mathcal{P}^{i}$-to-$\mathcal{R}^{\mathcal{N}}$ iteration to $\mathcal{W}$. Let $\mathcal{W}^{**}$ be the result of a fully backgrounded construction relative to $\Sigma_{\mathcal{R}^{\mathcal{N}}}$ inside $\mathcal{N}$. We can now apply universality of background constructions to the comparison of $\mathcal{W}^{*}$ and $\mathcal{W}^{**}$ in $\mathcal{W}$ and finish the proof as in [13, Lemma 2.16]. ∎
@@
As in [13, Lemma 2.19] the operator $S_{i}^{\mathcal{Q}_{i}}$ does not depend on the choice of $\mathcal{Q}_{i}$, so we can simplify the notation and write $S_{i}(X)$ for $S_{i}^{\mathcal{Q}_{i}}(X)$ for some $\kappa_{i}$-good universal mouse $\mathcal{Q}_{i}$ over $X$.
@@
Let $i<\omega$. We say an $\mathcal{M}|\delta_{i+1}$-extender $E$ with critical point $\kappa_{i}$ is characterized via $S_{i}$ iff
$Hull_{1}^{S_{i}(\mathcal{M}|\xi)}(\mathcal{M}|\kappa_{i})=S_{i}(\mathcal{M}|\kappa_{i})$
where $\xi\geq\kappa_{i}^{+}$ is the least $i$-weak cutpoint [^18] Recall that an ordinal $\alpha$ is an $i$-weak cutpoint above $\kappa_{i}$ if all extenders overlapping $\alpha$ have critical point $\kappa_{j}$ for some $j<i$. of $\operatorname{Ult}(\mathcal{M},E)$ above $\kappa_{i}$, and if $\alpha$ is the successor of $\xi$ in $\operatorname{Ult}(\mathcal{M},E)$, then $E$ is the $(\kappa_{i},\alpha)$-extender derived from the uncollapse map
$k\colon S_{i}(\mathcal{M}|\kappa_{i})\rightarrow S_{i}(\mathcal{M}|\xi).$
@@
Let $i<\omega$. An $\mathcal{M}|\delta_{i+1}$-extender $E$ with critical point $\kappa_{i}$ is certified iff it is characterized via the operator $S_{i}$.

### v3

@@ 5.5. All extenders on the $\mathcal{M}$-sequence are certified
Let $i\geq 0$, suppose $(\mathsf{IH})_{i}$ and let $E$ be an extender on the sequence of $\mathcal{M}^{(i)}$ with critical point $\kappa_{i}$ that is added by the fully backgrounded part of the construction of $\mathcal{M}^{(i)}$. Then $E$ is certified in $\mathcal{W}$.
@@
Let $F$ be an extender with critical point $\kappa_{i}$ that gets added by the fully backgrounded part of the construction of $\mathcal{M}^{(i)}$. Say $\mathcal{N}_{\xi+1}=(J_{\alpha}^{\vec{E}},\in,\vec{E},F)$ and let $F^{*}$ be the background extender of $F$. Let $\mathcal{U}$ be an iteration tree on $\mathcal{P}_{\infty}^{i,-}$ of limit length such that if we let $b=\Sigma^{i,\infty}(\mathcal{U})$,
$\pi^{\Sigma^{i,\infty}}_{\mathcal{P}_{\infty}^{i,-},\pi_{F}(\mathcal{P}_{\infty}^{i,-})}=\pi_{b}^{\mathcal{U}}.$
Moreover, let $\mathcal{U}^{*}$ be an iteration tree on $\mathcal{P}_{\infty}^{i,-}$ of limit length such that if we let $b^{*}=\Sigma^{i,\infty}(\mathcal{U}^{*})$,
$\pi^{\Sigma^{i,\infty}}_{\mathcal{P}_{\infty}^{i,-},\pi_{F^{*}}(\mathcal{P}_{\infty}^{i,-})}=\pi_{b^{*}}^{\mathcal{U}^{*}}.$
@@
be the canonical realization map. Then $k\upharpoonright\pi_{F}(\mathcal{P}_{\infty}^{i,-})\colon\pi_{F}(\mathcal{P}_{\infty}^{i,-})\rightarrow\pi_{F^{*}}(\mathcal{P}_{\infty}^{i,-})$ makes the diagram in Figure 2 commute.
$\pi_{F^{*}}$$\pi_{F^{*}}(\mathcal{P}_{\infty}^{i,-})$$\mathcal{P}_{\infty}^{i,-}$$\pi_{F}$$\pi_{F}(\mathcal{P}_{\infty}^{i,-})$$k$
@@
As $F^{*}$ is an extender on the sequence of $\mathcal{W}$ and $(\Sigma_{i})^{g}$ is strongly guided by term relations for all $g\subseteq\operatorname{Col}(\omega,\mathcal{P}^{i})$,
$\pi_{F^{*}}\upharpoonright\mathcal{P}_{\infty}^{i,-}=\pi_{b^{*}}^{\mathcal{U}^{*}}.$
We now define a cofinal branch $c$ through $\mathcal{U}$ such that $k(\mathcal{M}_{c}^{\mathcal{U}})=\mathcal{M}_{b^{*}}^{\mathcal{U}^{*}}$ and $\pi_{F}\upharpoonright\mathcal{P}_{\infty}^{i,-}=\pi_{c}^{\mathcal{U}}$.
@@
Both $\mathcal{U}^{\frown}c$ and $\mathcal{U}^{*}{}^{\frown}b^{*}$ are iteration trees on $\mathcal{P}_{\infty}^{i,-}$ and $\mathcal{U}^{\frown}c$ is a hull of $\mathcal{U}^{*}{}^{\frown}b^{*}$. As $\mathcal{U}^{*}{}^{\frown}b^{*}$ is according to $\Sigma^{i,\infty}$, by hull condensation $\mathcal{U}^{\frown}c$ is according to $\Sigma^{i,\infty}$ as well and hence
@@
Let $i\geq 0$, suppose $(\mathsf{IH})_{i}$ and let $E$ be an extender on the sequence of $\mathcal{M}$ with critical point $\kappa_{i}$ that is added by the generically countably complete part of the construction of $\mathcal{M}$ [^26] We are not assuming that $\mathcal{M}$ is a proper class model here. The construction of $\mathcal{M}$ might break down at some stage after $E$ was added. Then $E$ is certified in $\mathcal{W}$.
@@
Let $E$ be an extender on the sequence of $\mathcal{M}$ with critical point $\kappa_{i}$ that is added by the generically countably complete part of the construction of $\mathcal{M}$. Let $\hat{\mathcal{W}}$ be an elementary substructure of size ${<}\kappa_{i}$ of a sufficiently large initial segment $\mathcal{W}|\Omega$ of $\mathcal{W}$ with $\mathcal{W}|(\delta_{i}^{+\omega})^{\mathcal{W}}\subseteq\hat{\mathcal{W}}$. Let
@@
and suppose that all relevant objects are in the range of $\sigma$. For notational simplicity write $\mathcal{M}$ for the result of the $\mathcal{M}$-construction in $\mathcal{W}|\Omega$ and let $\hat{\mathcal{M}}$ be the transitive collapse of $\mathcal{M}$, i.e., the result of the $\mathcal{M}$-construction in $\hat{\mathcal{W}}$. Say $\sigma(\hat{E},\hat{\mathcal{P}}_{\infty}^{i},\hat{\Sigma}^{i,\infty})=(E,\mathcal{P}_{\infty}^{i},\Sigma^{i,\infty})$.
Then, by hull condensation, $\hat{\mathcal{P}}_{\infty}^{i}$ is an iterate of $\mathcal{P}^{i}$ via $\Sigma^{i}$. Hence, $\hat{\mathcal{P}}_{\infty}^{i}$ is an element of the direct limit system giving rise to $\mathcal{P}_{\infty}^{i}$. So when comparing $\hat{\mathcal{P}}_{\infty}^{i}$ via $\hat{\Sigma}^{i,\infty}$ and $\mathcal{P}_{\infty}^{i}$ via $\Sigma^{i,\infty}$ inside $\mathcal{W}$, $\hat{\mathcal{P}}_{\infty}^{i}$ iterates to $\mathcal{P}_{\infty}^{i}$. Moreover,
$\sigma\upharpoonright\hat{\mathcal{P}}_{\infty}^{i,-}=\pi^{\hat{\Sigma}^{i,\infty}}_{\hat{\mathcal{P}}_{\infty}^{i,-},\sigma(\hat{\mathcal{P}}_{\infty}^{i,-})}$
@@
$\pi_{\hat{E}}\upharpoonright\hat{\mathcal{P}}_{\infty}^{i,-}=\pi^{\hat{\Sigma}^{i,\infty}}_{\hat{\mathcal{P}}_{\infty}^{i,-},\pi_{\hat{E}}(\hat{\mathcal{P}}_{\infty}^{i,-})}.$
So $\hat{E}$ is certified via $\hat{\Sigma}^{i,\infty}$. By elementarity of $\sigma$, this implies that $E$ is certified via $\Sigma^{i,\infty}$, as desired. ∎
@@ 5.6. Characterization via operators
Now we argue that we can also characterize extenders using operators in order to obtain the extenders in a simply definable way (e.g., in $M_{1}^{\#,\Sigma_{i}}$ of the current level). In what follows, we define universal mice at $\kappa_{i}$ and $\kappa_{i}$-good universal mice as in [14]. The arguments in [14] generalize to $\kappa_{i}$, $i>0$, by considering $\mathcal{M}$-constructions instead of fully backgrounded constructions. In particular, [14, Lemmas 2.15 and 2.16] generalize and we can show that good universal mice exist. We give the details in what follows.
@@
Let $i<\omega$, suppose $(\mathsf{IH})_{i}$ and let $X\in\mathcal{W}|\delta_{i+1}$ be a premouse such that $\mathcal{M}^{(i)}|\delta_{i}\unlhd X$. We say $\mathcal{Q}\subseteq\mathcal{W}|\delta_{i+1}$ is a universal mouse over $X$ at $\kappa_{i}$ if $\mathcal{Q}$ is an $X$-premouse such that $\mathcal{Q}$ is $(\delta_{i+1},\delta_{i+1})$-iterable with respect to the extenders on its sequence with critical point above $X\cap\operatorname{Ord}$, for every strong cutpoint $\eta\in(X\cap\operatorname{Ord},\delta_{i+1})$ in $\mathcal{Q}$, $\mathcal{O}_{\eta}^{\mathcal{Q}}=Lp^{\infty}(\mathcal{Q}|\eta)$ and for a stationary set of $\xi<\delta_{i+1}$, $(\xi^{+})^{\mathcal{Q}}=(\xi^{+})^{\mathcal{W}}$. Moreover, in case $i>0$, $\mathcal{Q}$ is $\kappa_{j}$-good for all $j<i$, i.e., whenever $\mathcal{R}^{\mathcal{Q}}$ (defined inductively as described below) is a $\Sigma_{j}$-iterate of $\mathcal{P}^{j}$ and whenever $g$ is generic over $\mathcal{W}$ for a partial order in $\mathcal{Q}$, $(\Sigma_{j}^{g})_{\mathcal{R}^{\mathcal{Q}}}\upharpoonright\mathcal{Q}[g]\in L_{1}(\mathcal{Q}[g])$.
Before we can define good universal mice (and the object $\mathcal{R}^{\mathcal{Q}}$ used in the previous definition), we need some additional notation. Let $i<\omega$, suppose $(\mathsf{IH})_{i}$ and let $X\in\mathcal{W}|\delta_{i+1}$ be a premouse with $X\unrhd\mathcal{M}^{(i)}|\delta_{i}$. For any $X$-premouse $\mathcal{Q}\subseteq\mathcal{W}|\delta_{i+1}$, let $\kappa^{\mathcal{Q}}$ denote the least ${<}\delta_{i+1}$-strong cardinal of $\mathcal{Q}$, if it exists. The reader might wish to recall the definition of being $\kappa_{i}$-suitable (see Definition 3.7) at this point. For any universal mouse $\mathcal{Q}\subseteq\mathcal{W}|\delta_{i+1}$ over $X$ at $\kappa_{i}$, consider, analogous to Section 5.2, the directed system given by
@@
Note that for an $\mathcal{R}$ as in the definition of $\mathcal{I}^{(i)}_{\mathcal{Q}}$ the notion of $\kappa_{i}$-suitability is well-defined inside $\mathcal{Q}$ as $\mathcal{Q}$ is sufficiently closed under the operation $X\mapsto Lp^{\Sigma_{\mathcal{R}},\infty}(X)$ by the proof of Lemma 5.9. Write $\mathcal{P}_{\infty}^{\mathcal{Q}}$ for the corresponding direct limit and $\delta_{\infty}^{\mathcal{Q}}$ for the largest Woodin cardinal in $\mathcal{P}_{\infty}^{\mathcal{Q}}$. Note that $\mathcal{P}_{\infty}^{\mathcal{Q}}\in\mathcal{Q}$ and let
$\mathcal{R}^{\mathcal{Q}}=\mathcal{P}_{\infty}^{\mathcal{Q}}|((\delta_{\infty}^{\mathcal{Q}})^{+\omega})^{\mathcal{P}_{\infty}^{\mathcal{Q}}}.$
@@
Let $i<\omega$, suppose $(\mathsf{IH})_{i}$ and let $X\in\mathcal{W}|\delta_{i+1}$ be a premouse such that $\mathcal{M}^{(i)}|\delta_{i}\unlhd X$. We say a universal mouse $\mathcal{Q}$ over $X$ at $\kappa_{i}$ is a $\kappa_{i}$-good universal mouse over $X$ iff it has a $(\delta_{i+1},\delta_{i+1})$-iteration strategy $\Phi$ with respect to extenders with critical point above $X\cap\operatorname{Ord}$ such that whenever $i\colon\mathcal{Q}\rightarrow\mathcal{N}$ is an embedding obtained from an iteration according to $\Phi$, $\mathcal{N}$ is a universal mouse, $\mathcal{R}^{\mathcal{N}}$ is a $\Sigma_{i}$-iterate of $\mathcal{P}^{i}$, and whenever $g$ is generic over $\mathcal{W}$ for a partial order in $\mathcal{N}$,
@@
The existence of good universal mice follows as in [14, Lemmas 2.15 and 2.16], cf. also the proof of Lemma 5.12.
@@
Let $i<\omega$, suppose $(\mathsf{IH})_{i}$ and let $X\unrhd\mathcal{M}^{(i)}|\delta_{i}$ be a premouse with $X\in\mathcal{W}|\delta_{i+1}$. Fix some ordinal $\eta$ with $X\cap\operatorname{Ord}<\eta<\delta_{i+1}$ and let $\mathcal{Q}=\mathcal{Q}^{X,\eta}$ be the result of an $\mathcal{M}$-construction above $X$ inside $\mathcal{W}|\delta_{i+1}$ where the fully backgrounded part only uses extenders with critical point above $\eta$. Let $\Phi$ be the iteration strategy for $\mathcal{Q}$ induced by the background strategy [^27] That means, $\Phi$ only acts on the extenders on the sequence of $\mathcal{Q}$ that are backgrounded. and let $\mathcal{N}$ be an iterate of $\mathcal{Q}$ (above $X$) according to $\Phi$ such that the iteration embedding $i\colon\mathcal{Q}\rightarrow\mathcal{N}$ exists. Then $\mathcal{N}$ is a $\kappa_{i}$-good universal mouse over $X$.
@@
$\mathcal{R}^{\mathcal{N}}$ is a $\Sigma_{i}$-iterate of $\mathcal{P}^{i}$ and $\mathcal{N}$ is $\kappa_{i}$-good, i.e., whenever $g$ is generic over $\mathcal{W}$ for a partial order in $\mathcal{N}$, $(\Sigma^{g}_{i})_{\mathcal{R}^{\mathcal{N}}}\upharpoonright\mathcal{N}[g]\in L_{1}(\mathcal{N}[g])$, by the argument for Lemma 5.12. So we only sketch the argument that $\mathcal{N}$ is universal and show that for a stationary set of $\xi<\delta_{i+1}$, $(\xi^{+})^{\mathcal{N}}=(\xi^{+})^{\mathcal{W}}$. Let $\mathcal{W}^{*}$ be the iterate of $\mathcal{W}$ obtained by applying the $\mathcal{P}^{i}$-to-$\mathcal{R}^{\mathcal{N}}$ iteration to $\mathcal{W}$. Let $\mathcal{W}^{**}$ be the result of a fully backgrounded construction relative to $\Sigma_{\mathcal{R}^{\mathcal{N}}}$ inside $\mathcal{N}$. We can now apply universality of background constructions to the comparison of $\mathcal{W}^{*}$ and $\mathcal{W}^{**}$ in $\mathcal{W}$ and finish the proof as in [14, Lemma 2.16]. ∎
@@
As in [14, Lemma 2.19] the operator $S_{i}^{\mathcal{Q}_{i}}$ does not depend on the choice of $\mathcal{Q}_{i}$, so we can simplify the notation and write $S_{i}(X)$ for $S_{i}^{\mathcal{Q}_{i}}(X)$ for some $\kappa_{i}$-good universal mouse $\mathcal{Q}_{i}$ over $X$.
@@
Let $i<\omega$ and suppose $(\mathsf{IH})_{i}$. We say an $\mathcal{M}^{(i)}|\delta_{i+1}$-extender $E$ with critical point $\kappa_{i}$ is characterized via $S_{i}$ iff [^28] Note that $\mathcal{M}^{(i)}|\xi=\operatorname{Ult}(\mathcal{M}^{(i)},E)|\xi$.
$Hull_{1}^{S_{i}(\mathcal{M}^{(i)}|\xi)}(\mathcal{M}^{(i)}|\kappa_{i})=S_{i}(\mathcal{M}^{(i)}|\kappa_{i})$
where $\xi\geq\kappa_{i}^{+}$ is the least $i$-weak cutpoint [^29] Recall that an ordinal $\alpha$ is an $i$-weak cutpoint above $\kappa_{i}$ if all extenders overlapping $\alpha$ have critical point $\kappa_{j}$ for some $j<i$. of $\operatorname{Ult}(\mathcal{M}^{(i)},E)$ above $\kappa_{i}$, and if $\alpha$ is the successor of $\xi$ in $\operatorname{Ult}(\mathcal{M}^{(i)},E)$, then $E$ is the $(\kappa_{i},\alpha)$-extender [^30] Recall that by our indexing convention, $E$ with critical point $\kappa_{i}$ will be indexed at $\alpha$. derived from the uncollapse map
$k\colon S_{i}(\mathcal{M}^{(i)}|\kappa_{i})\rightarrow S_{i}(\mathcal{M}^{(i)}|\xi).$
@@
Let $i<\omega$ and suppose $(\mathsf{IH})_{i}$. An $\mathcal{M}^{(i)}|\delta_{i+1}$-extender $E$ with critical point $\kappa_{i}$ is certified iff it is characterized via the operator $S_{i}$.